OpenStudy (anonymous):

what multiplies to 2x^2 and adds to -1x?

OpenStudy (anonymous):

im factoring, the problem is 2x^2 -x-1

jimthompson5910 (jim_thompson5910):

multiply first and last coefficients: 2*(-1) = -2 find two numbers that multiply to -2 AND add to -1 (the middle term) these two numbers are -2 and 1 since -2*1 = -2 and -2+1 = -1

jimthompson5910 (jim_thompson5910):

so this means 2x^2-x-1 2x^2+x-2x-1 (2x^2+x)+(-2x-1) x(2x+1)+(-2x-1) x(2x+1)-1(2x+1) I'll let you finish

OpenStudy (anonymous):

sweeet! thank you!

jimthompson5910 (jim_thompson5910):

tell me what you get

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

(x-1) (2x+1)

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yea, im tired couldnt think of the numbers tonight. thank you!

jimthompson5910 (jim_thompson5910):

sure thing